# Spatial Disparities in Fire Shelter Accessibility: Capacity Challenges in the Palisades and Eaton Fires

### v1

@@ 4.4 Workflow of This Study
    Case 4.2: Distance-based Approach — The intuition behind this approach is to provide evacuees with the nearest possible shelter locations, thereby reducing the time and distance needed to reach shelters. Our distance-based greedy algorithm is described as follows. Starting from the demand areas, the algorithm performs a buffer-based search, incrementally increasing the buffer distance (for example, from 1 mile to 2 miles, and so on). The algorithm continues expanding buffer distances until the total shelter capacity reaches the required demand.
@@
Case 2 I3. 8 shelters opened on Jan 12th I8. Population within evacuation order & warning zones, excluding areas affected by active fires I9. Road network excluding roads within fire perimeters Not considered P6. Compute accessibility using specified supply, demand, and road network Figures 6 & 8: Accessibility with road closures
@@
Case 4 Varied hypothetical shelters I6. All population in evacuation order & warning zones I4. Full road network I7. Considered P8. Compute accessibility using specified supply, demand, and road network with congestion Figures 9 & 10: Accessibility based on hypothetical shelters sized for all residents Figure 11: Accessibility to 8 shelters opened on Jan 12th
@@ 5.1 Distance to nearest shelters
General insights into travel times to shelters can be drawn from Figure 5. In this map, the color shading of each grid cell represents estimated travel times to the nearest shelters under conditions without road congestion. For the Palisades Fire, the southeastern regions within the evacuation order and warning zones generally have shelters located within approximately 3 to 8 minutes, indicating relatively easy access compared to other areas. In contrast, travel times from the inner mountainous regions can extend up to 40 minutes. For the Eaton Fire, the Pasadena Civic Auditorium is the closest shelter for residents in the evacuation order zones. Travel times to this location are generally up to 20 minutes, except for the mountainous areas. In these sparsely populated mountain regions, travel times are more than double those in residential zones. Figure 5 illustrates travel times only to the nearest shelters whose capacities were significantly smaller than the demand as discussed in the section 4.4; as a result, many evacuees had to travel to shelters farther away than the closest ones.
@@ 5.2 Shelter accessibility under congested and uncongested traffic conditions
Figure 6: Shelter accessibility from evacuation order and warning zones on January 12, assuming road closures but no traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-6
Figures 7 and 8 build upon this baseline to examine the impact of traffic congestion on spatial accessibility. Figure 7 reflects congested traffic conditions, as defined in Case 2 of Table 2 and Figure 4, where travel times were significantly increased due to congestion. In this scenario, not only were roads within the fire perimeters closed, but travel within a five-kilometer buffer zone around the evacuation areas was assumed to be extremely limited, with speeds reduced to just 10 kph.
Figure 7: Shelter accessibility from evacuation order and warning zones on January 12, accounting for road closures and traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-7
@@
Figure 8: Shelter accessibility from evacuation order and warning zones on January 12, accounting for road closures but without traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-8
The comparison between Figures 7 and 8 highlights the negative impact of traffic congestion on equitable access to shelters. Figure 8, which depicts accessibility without congestion, shows generally high access across most of the study area–except near active fire perimeters and in regions with sparse road networks, particularly in the southwestern Palisades zones and the mountainous areas of the Eaton evacuation zones. In contrast, Figure 7, which incorporates the effects of traffic congestion, reveals a substantial decline in accessibility. Only a few areas–such as narrow areas the northwest and southeast Palisades fire evacuated zones and a small region near the Pasadena Civic Auditorium near Eaton fire evacuated zones—maintain high accessibility scores. Most other areas fall into the Very Poor to Moderate accessible categories, indicating a widespread decline in shelter access under congested conditions.
This disparity is further supported by Gini coefficients, which quantify accessibility inequality. The coefficients for Figures 7 and 8 are 0.45 and 0.09, respectively, demonstrating a significant increase in inequality under congested traffic conditions.
@@ 5.3 Shelter accessibility with hypothetical shelter locations
As discussed in Section 4.4, approximately 81,387 individuals were left without access to designated shelters. In this section, we explore strategies for the placement of additional shelters to accommodate the entire at-risk population and improve overall accessibility during wildfire emergencies. Specifically, we apply two approaches introduced in Section 4.4–the Capacity-based Approach (Case 4.1) and the Distance-based Approach (Case 4.2). Both cases assume traffic congestion, with travel speeds reduced to 10 kph within a five-kilometer buffer zone surrounding the evacuation areas, consistent with the previous scenario.
Figure 9 presents the results of the Capacity-based Approach, as outlined in Case 4.1 of Section 4.4. In this scenario, the first eleven shelters used during the Palisades and Eaton wildfires were included, providing a combined capacity for 6,424 evacuees–significantly below the total at-risk population of 86,661 within the evacuation zones. To address this shortfall, the algorithm began from demand areas (i.e., the Palisades and Eaton zones) and incrementally searched for additional potential shelters drawn from the National Shelter System Facilities dataset ([51]) by expanding buffer distances outward. Shelters were added one by one until the combined capacity reached twice the estimated demand (173,222 people) to ensure sufficient coverage. This threshold was reached at approximately 13 kilometers from the evacuation zones. At that point, shelter selection was finalized by choosing facilities within the 13-kilometer buffer in descending order of capacity–starting with the largest–until the cumulative shelter capacity met or exceeded the population in need (86,661 people). Through this process, the selected shelters shown in Figure 9 were identified, offering a total combined capacity of 86,957 evacuees.
Figure 9: Shelter accessibility based on hypothetical shelters sized to accommodate all residents. A greedy algorithm prioritized shelters with the largest capacities until population needs were met. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-9
@@
Figure 10: Shelter accessibility based on hypothetical shelters sized to accommodate all residents. Shelters were incrementally added in order of proximity until the entire population in evacuation and warning zones was covered. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-10
Shelter placements using the Capacity-based Approach (Figure 9) and the Distance-based Approach (Figure 10) show greater improvements in accessibility compared to the actual eleven shelters opened on January 12, as shown in Figure 11. Figure 11 presents the accessibility scores under the same traffic congestion scenario depicted in Figures 9 and 10. To ensure a fair comparison, the color scheme and class intervals in Figure 11 were aligned with those in Figures 9 and 10. As expected, the overall accessibility scores in the hypothetical shelter scenarios (Figures 9 and 10) are generally higher across the study area due to the increased number of shelters, compared to the case of eleven shelters in Figure 11. When examining equity in shelter access across the three scenarios, the Gini coefficient is 0.34 for the Capacity-based Approach and 0.31 for the Distance-based Approach (Figure 10), while the original 11-shelter scenario (Figure 11) shows a much higher Gini coefficient of 0.69. This indicates significantly greater disparities in shelter access under the actual 11-shelter scenario.
Between the two hypothetical strategies in Figure 9 and 10, the Distance-based Approach shows slightly better equity, as indicated by its lower Gini coefficient. This suggests that the Distance-based approach provides more equal access to shelters than the Capacity-based Approach. The likely reason is that it prioritizes the minimum travel distance and time for evacuees, while the Capacity-based Algorithm focuses on reducing the number of shelters required, emphasizing efficiency and logistical resource distribution.
Figure 11: Shelter accessibility from evacuation order and warning zones on January 12, assuming full road network availability and traffic congestion. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-11
@@ Appendix A Shelter Estimation
    Shelter location:Van Nuys/Sherman Oaks Recreation Center
    Estimating the shelter capacity for the Van Nuys/Sherman Oaks Recreation Center at 14201 Huston St., Sherman Oaks, CA, involves assessing the available indoor space. The recreation center features a 15,000-square-foot gymnasium, which serves as a significant indoor space suitable for sheltering purposes [52]. Standard estimates suggest that about 30% of the total indoor space is unusable due to hallways, storage areas, restrooms, and other non-shelterable spaces. With a total indoor space of 15,000 square feet, the estimated usable space is 10,500 square feet, calculated as 15,000 sq. ft. $\times$ 0.70. Based on an assumption that each person requires 100 square feet, the estimated shelter capacity for the facility is approximately 105 people.
@@
    Shelter location: Calvary Community Church at 5495 Via Rocas, Westlake Village, CA
    The building area measured from the satellite image is 10,573.27 $m^{2}$. After excluding 30% for unusable space, the estimated shelter capacity for Calvary Community Church is approximately 797 people, assuming that each person requires 100 sq. ft.
@@
    Shelter location: Pasadena Civic Auditorium
    Estimating the shelter capacity of the Pasadena Convention Center at 300 E Green Street, Pasadena, CA, involves determining the usable indoor area. The convention center offers approximately 130,000 square feet of flexible event space, including exhibit halls, ballrooms, and meeting rooms [75]. Without detailed floor plans, it is standard to estimate that about 30% of the total space is unusable due to hallways, storage areas, restrooms, and other non-shelterable areas. Given a total building area of 130,000 square feet, the estimated unusable space is 39,000 square feet, resulting in a usable area of approximately 91,000 square feet. Assuming that each person requires 100 square feet, the estimated shelter capacity for the facility is approximately 105 people.
@@
    Shelter location: Pomona Fairplex
    The Fairplex offers approximately 325,000 square feet of indoor exhibit space, with several major exhibition halls that can be utilized for sheltering purposes [76]. Building 4, the largest of these spaces, covers approximately 105,600 square feet with a length of 770 feet and a width ranging from 130 to 146 feet. It also features a stage measuring 64 by 19 feet and includes eight dressing rooms. Buildings 5, 6, 7, and 8 each provide 33,600 square feet, with dimensions of 100 feet by 318 feet. Each of these buildings is equipped with its own ticket office and two 360-square-foot show support offices. Building 3, covering approximately 11,486 square feet, is divided into five interior rooms, offering flexibility for various uses. These facilities are designed to accommodate a wide range of events, from large-scale exhibitions to more intimate gatherings [76]. The total indoor area of the major exhibition halls, including Building 4 (105,600 sq. ft.), Buildings 5–8 (33,600 sq. ft. each), and Building 3 (11,486 sq. ft.), sums to approximately 150,686 square feet. After excluding 30% for unusable space, the estimated shelter capacity for the facility is 1,056 people, assuming that each person requires 100 square feet.
@@
    Shelter location: Glendale Civic Center
    Estimating the shelter capacity for the Glendale Civic Center at 613 E Broadway, Glendale, CA 91206, involves assessing the available indoor areas and applying FEMA and Red Cross Emergency Shelter Standards. The Glendale Civic Center provides a variety of spaces suitable for sheltering purposes, with a total meeting space of approximately 28,985 square feet [77]. Considering that 30% of the space is typically unusable due to hallways, storage, and other non-shelterable areas, the estimated usable space is approximately 20,290 square feet. Based on the assumption that each person requires 100 square feet, the estimated shelter capacity for the facility is 202 people .

### v2

@@ 18    $C_{\text{refined}}\leftarrow C_{\text{refined}}+C(s)$;
— The intuition behind this approach is to provide evacuees with the nearest possible shelter locations, thereby reducing the time and distance needed to reach shelters. Our distance-based greedy algorithm is described as follows. Starting from the demand areas, the algorithm performs a buffer-based search, incrementally increasing the buffer distance (for example, from 1 mile to 2 miles, and so on). The algorithm continues expanding buffer distances until the total shelter capacity reaches the required demand.
Algorithm 2 presents the distance-based approach in pseudo code. The algorithm initializes an empty set to hold the final selection of shelters (Line 1). For each demand area (Line 2), it begins with an initial buffer radius $\delta$ (Line 5) and an empty candidate set (Line 3), as well as a capacity counter (Line 4). Within the while-loop (Lines 6–11), the algorithm identifies all shelters within the current radius (Line 7). Newly discovered shelters are added to the candidate list, and their capacities are accumulated (Lines 8–9). If the current total capacity is still below the required demand, the buffer is increased (Line 10), and the process repeats. Once the cumulative capacity meets or exceeds the demand, all candidate shelters found within the final buffer are added to the overall selection (Line 12). The algorithm proceeds to the next demand area and finally returns the complete list of selected shelters (Line 13).
Input: Set of demand areas $D$
Input: Set of shelters $S$ with capacities $C(s)$ and locations
Input: Population demand $P_{d}$ for each $d\in D$
Input: Initial buffer distance $\delta$
Input: Buffer increment $\Delta$
Output: Set of selected shelters $S_{\text{selected}}$
1
$S_{\text{selected}}\leftarrow\emptyset$ ;
// Initialize the final list of selected shelters
2
3foreach demand area $d\in D$ do
 $S_{\text{candidate}}\leftarrow\emptyset$ ;
 // Temporary set of nearby shelters
 $C_{\text{total}}\leftarrow 0$ ;
 // Total capacity found so far
 $r\leftarrow\delta$ ;
 // Start with initial buffer radius
4
5 while $C_{\text{total}}<P_{d}$ do
6      Find shelters $S_{r}\subseteq S$ within radius $r$ of demand area $d$;
7    foreach $s\in S_{r}\setminus S_{\text{candidate}}$ do
8         Add $s$ to $S_{\text{candidate}}$;
9       $C_{\text{total}}\leftarrow C_{\text{total}}+C(s)$;
10
11    if $C_{\text{total}}<P_{d}$ then
       $r\leftarrow r+\Delta$ ;
       // Expand buffer and continue search
12
13
14
15  Add all shelters in $S_{\text{candidate}}$ to $S_{\text{selected}}$;
16
17
return $S_{\text{selected}}$ ;
// Return shelters selected for all demand areas
18
Algorithm 2 Distance-Based Shelter Selection
@@ 15  Add all shelters in $S_{\text{candidate}}$ to $S_{\text{selected}}$;
Case 2 I3. 8 shelters opened on Jan 12th I8. Population within evacuation order & warning zones, excluding areas affected by active fires I9. Road network excluding roads within fire perimeters Not considered P6. Compute accessibility using specified supply, demand, and road network Figures 5 & 7: Accessibility with road closures
@@
Case 4 Varied hypothetical shelters I6. All population in evacuation order & warning zones I4. Full road network I7. Considered P8. Compute accessibility using specified supply, demand, and road network with congestion Figures 8 & 9: Accessibility based on hypothetical shelters sized for all residents Figure 10: Accessibility to 8 shelters opened on Jan 12th
@@ 5.1 Distance to nearest shelters
General insights into travel times to shelters can be drawn from Figure 4. In this map, the color shading of each grid cell represents estimated travel times to the nearest shelters under conditions without road congestion. For the Palisades Fire, the southeastern regions within the evacuation order and warning zones generally have shelters located within approximately 3 to 8 minutes, indicating relatively easy access compared to other areas. In contrast, travel times from the inner mountainous regions can extend up to 40 minutes. For the Eaton Fire, the Pasadena Civic Auditorium is the closest shelter for residents in the evacuation order zones. Travel times to this location are generally up to 20 minutes, except for the mountainous areas. In these sparsely populated mountain regions, travel times are more than double those in residential zones. Figure 4 illustrates travel times only to the nearest shelters whose capacities were significantly smaller than the demand as discussed in Section D; as a result, many evacuees had to travel to shelters farther away than the closest ones.
@@ 5.2 Shelter accessibility under congested and uncongested traffic conditions
Figure 5: Shelter accessibility from evacuation order and warning zones on January 12, assuming road closures but no traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-5
Figures 6 and 7 build upon this baseline to examine the impact of traffic congestion on spatial accessibility. Figure 6 reflects congested traffic conditions, as defined in Case 2 of Table 2 and Figure 3, where travel times were significantly increased due to congestion. In this scenario, not only were roads within the fire perimeters closed, but travel within a five-kilometer buffer zone around the evacuation areas was assumed to be extremely limited, with speeds reduced to just 10 kph.
Figure 6: Shelter accessibility from evacuation order and warning zones on January 12, accounting for road closures and traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-6
@@
Figure 7: Shelter accessibility from evacuation order and warning zones on January 12, accounting for road closures but without traffic congestion. Interactive version available at: https://sites.google.com/view/palisades-and-eton-fires/figure-7
The comparison between Figures 6 and 7 highlights the negative impact of traffic congestion on equitable access to shelters. Figure 7, which depicts accessibility without congestion, shows generally high access across most of the study area–except near active fire perimeters and in regions with sparse road networks, particularly in the southwestern Palisades zones and the mountainous areas of the Eaton evacuation zones. In contrast, Figure 6, which incorporates the effects of traffic congestion, reveals a substantial decline in accessibility. Only a few areas–such as narrow areas the northwest and southeast Palisades fire evacuated zones and a small region near the Pasadena Civic Auditorium near Eaton fire evacuated zones—maintain high accessibility scores. Most other areas fall into the Very Poor to Moderate accessible categories, indicating a widespread decline in shelter access under congested conditions.
This disparity is further supported by Gini coefficients, which quantify accessibility inequality. The coefficients for Figures 6 and 7 are 0.45 and 0.09, respectively, demonstrating a significant increase in inequality under congested traffic conditions.
@@ 5.3 Shelter accessibility with hypothetical shelter locations
As discussed in Section D, approximately 81,387 individuals were left without access to designated shelters. In this section, we explore strategies for the placement of additional shelters to accommodate the entire at-risk population and improve overall accessibility during wildfire emergencies. Specifically, we apply two approaches introduced in Section D–the Capacity-based Approach (Case 4.1) and the Distance-based Approach (Case 4.2). Both cases assume traffic congestion, with travel speeds reduced to 10 kph within a five-kilometer buffer zone surrounding the evacuation areas, consistent with the previous scenario.
Figure 8 presents the results of the Capacity-based Approach, as outlined in Case 4.1 of Section D. In this scenario, the first eleven shelters used during the Palisades and Eaton wildfires were included, providing a combined capacity for 6,424 evacuees–significantly below the total at-risk population of 86,661 within the evacuation zones. To address this shortfall, the algorithm began from demand areas (i.e., the Palisades and Eaton zones) and incrementally searched for additional potential shelters drawn from the National Shelter System Facilities dataset ([35]) by expanding buffer distances outward. Shelters were added one by one until the combined capacity reached twice the estimated demand (173,222 people) to ensure sufficient coverage. This threshold was reached at approximately 13 kilometers from the evacuation zones. At that point, shelter selection was finalized by choosing facilities within the 13-kilometer buffer in descending order of capacity–starting with the largest–until the cumulative shelter capacity met or exceeded the population in need (86,661 people). Through this process, the selected shelters shown in Figure 8 were identified, offering a total combined capacity of 86,957 evacuees.
Figure 8: Shelter accessibility based on hypothetical shelters sized to accommodate all residents. A greedy algorithm prioritized shelters with the largest capacities until population needs were met. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-8
@@
Figure 9: Shelter accessibility based on hypothetical shelters sized to accommodate all residents. Shelters were incrementally added in order of proximity until the entire population in evacuation and warning zones was covered. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-9
Shelter placements using the Capacity-based Approach (Figure 8) and the Distance-based Approach (Figure 9) show greater improvements in accessibility compared to the actual eleven shelters opened on January 12, as shown in Figure 10. Figure 10 presents the accessibility scores under the same traffic congestion scenario depicted in Figures 8 and 9. To ensure a fair comparison, the color scheme and class intervals in Figure 10 were aligned with those in Figures 8 and 9. As expected, the overall accessibility scores in the hypothetical shelter scenarios (Figures 8 and 9) are generally higher across the study area due to the increased number of shelters, compared to the case of eleven shelters in Figure 10. When examining equity in shelter access across the three scenarios, the Gini coefficient is 0.34 for the Capacity-based Approach and 0.31 for the Distance-based Approach (Figure 9), while the original 11-shelter scenario (Figure 10) shows a much higher Gini coefficient of 0.69. This indicates significantly greater disparities in shelter access under the actual 11-shelter scenario.
Between the two hypothetical strategies in Figure 8 and 9, the Distance-based Approach shows slightly better equity, as indicated by its lower Gini coefficient. This suggests that the Distance-based approach provides more equal access to shelters than the Capacity-based Approach. The likely reason is that it prioritizes the minimum travel distance and time for evacuees, while the Capacity-based Algorithm focuses on reducing the number of shelters required, emphasizing efficiency and logistical resource distribution.
Figure 10: Shelter accessibility from evacuation order and warning zones on January 12, assuming full road network availability and traffic congestion. Interactive version: https://sites.google.com/view/palisades-and-eton-fires/figure-10
@@ Appendix A Shelter Estimation
This section provides detailed descriptions of how shelter capacities were estimated for major emergency shelters in the study area. Estimations were based on publicly available building size data, architectural plans, and FEMA/Red Cross shelter guidelines. A common assumption was that approximately 30% of total indoor space is unusable due to hallways, storage, restrooms, and other non-shelterable areas, and that each evacuee requires $100\text{\,}\mathrm{f}\mathrm{t}^{2}$ of space.
@@
    Van Nuys/Sherman Oaks Recreation Center
    Located at 14201 Huston St., Sherman Oaks, CA, the recreation center includes a gymnasium of approximately $15\,000\text{\,}\mathrm{f}\mathrm{t}^{2}$ [4]. Assuming 70% of the space is usable, the effective area is $10\,500\text{\,}\mathrm{f}\mathrm{t}^{2}$, yielding an estimated capacity of 105 people.
@@
    Calvary Community Church
    Situated at 5495 Via Rocas, Westlake Village, CA, the building covers approximately $10\,573.27\text{\,}\mathrm{m}^{2}$, as measured and verified using high-resolution satellite imagery. After excluding 30% unusable space, the usable area equals $7401.29\text{\,}\mathrm{m}^{2}$. Converting to $79\,720\text{\,}\mathrm{f}\mathrm{t}^{2}$, the estimated shelter capacity is 797 people.
@@
    Pasadena Civic Auditorium
    The Pasadena Convention Center (300 E Green Street, Pasadena, CA) provides roughly $130\,000\text{\,}\mathrm{f}\mathrm{t}^{2}$ of flexible indoor area [78]. Assuming 30% unusable space, the usable area is $91\,000\text{\,}\mathrm{f}\mathrm{t}^{2}$, supporting an estimated capacity of 910 people.
@@
    Pomona Fairplex
    The Fairplex complex offers about $325\,000\text{\,}\mathrm{f}\mathrm{t}^{2}$ of indoor exhibit space, including multiple halls suitable for emergency sheltering [15]. Building 4 ($105\,600\text{\,}\mathrm{f}\mathrm{t}^{2}$) and Buildings 5–8 ($3$ $3\,600\text{\,}\mathrm{f}\mathrm{t}^{2}$ each) plus Building 3 (11 $1\,486\text{\,}\mathrm{f}\mathrm{t}^{2}$) together total $150\,686\text{\,}\mathrm{f}\mathrm{t}^{2}$. Excluding 30% unusable space leaves $105\,480\text{\,}\mathrm{f}\mathrm{t}^{2}$, corresponding to an estimated capacity of 1,056 people.
@@
    Glendale Civic Center
    Located at 613 E Broadway, Glendale, CA 91206, the building area, measured from high-resolution satellite imagery, is approximately $28\,985\text{\,}\mathrm{f}\mathrm{t}^{2}$ ($2692.8\text{\,}\mathrm{m}^{2}$). Considering that about 30% of the space is typically unusable due to hallways, storage, and other non-shelterable areas, the estimated usable area is $20\,290\text{\,}\mathrm{f}\mathrm{t}^{2}$. Assuming an average allocation of $100\text{\,}\mathrm{f}\mathrm{t}^{2}$ per person, the estimated shelter capacity of the facility is approximately 202 people.
Table 3: Summary of estimated shelter capacities.
Shelter Name Usable Area (ft2) Assumed Area per Person (ft2) Estimated Capacity
Van Nuys/Sherman Oaks Recreation Center 10,500 100 105
Calvary Community Church 79,720 100 797
Pasadena Civic Auditorium 91,000 100 910
Pomona Fairplex 105,480 100 1,056
Glendale Civic Center 20,290 100 202
Appendix B Road Network
This section presents the visualization of the road network downloaded from OpenStreetMap (OSM) using the Python package OSMnx. Figure 11A shows both major and local roads across the entire study area, while Figures 11B and 11C highlight the Palisades and Eaton Fire areas. The results demonstrate sparse road density within forested zones inside wildfire perimeters, whereas populations are largely concentrated in regions with denser road networks.
Figure 11: Road network data retrieved from OpenStreetMap (OSM) for the study area. (A) Full regional road network; (B) zoomed view of the Palisades Fire area; (C) zoomed view of the Eaton Fire area.
Appendix C Travel-Time Thresholds and Distance-Decay Specification
In our spatial accessibility framework, we adopt a Gaussian distance-weighting function with a specified width ($\sigma=30$ minutes) and impose a catchment travel-time threshold of 120 minutes (2 hours) under free-flow conditions. The Gaussian kernel is designed such that the weight $W_{ij}$ decreases sharply as travel time increases, approaching near-zero around 90 minutes. This formulation captures the rapid decline in accessibility as the effort required to reach a shelter increases. The 120-minute threshold serves as a hard cutoff under uncongested travel speeds, meaning that any origin located more than two hours from the nearest shelter is considered outside the practical catchment (accessibility = 0).
This modeling strategy aligns with established practices in generalized and enhanced Two-Step Floating Catchment Area (2SFCA) methods, which require defining an upper distance or time limit beyond which access is considered negligible [47, 73]. Numerous studies have shown that accessibility results are highly sensitive to the choice of travel-time threshold. For example, [34] showed that reducing the maximum driving time from 60 to 30 minutes in Haiti decreased population coverage from 97.5% to 85.4%. Similarly, [24, 2] evaluated multiple travel-time thresholds (15, 30, and 60 minutes) and found that accessibility declines substantially under more restrictive assumptions. Building on these studies, our use of a 90-minute effective decay and a 120-minute cutoff provides an upper-bound estimate that avoids prematurely excluding remote or traffic-congested areas from the analysis.
Empirical evidence supports the plausibility of this assumption. The NIST Camp Fire case study documented evacuees trapped in vehicles for over 90 minutes during the 2018 wildfire due to gridlock and limited egress routes [14]. Similarly, FEMA’s all-hazard evacuation planning guidance recommends flexible clearance times that can exceed one to two hours depending on hazard type, population density, and network capacity [29]. These findings emphasize that while actual travel during evacuations may far exceed two hours, using a 2-hour free-flow threshold provides a consistent and conservative upper bound for analytical purposes. In summary, setting $\sigma=30$ minutes and a 120-minute travel-time threshold captures the steep decline in accessibility associated with longer travel while maintaining computational tractability and empirical validity. The 90-minute point represents an effective limit of meaningful access, while the 120-minute cutoff ensures that accessibility beyond realistic uncongested travel conditions is excluded. This approach integrates real-world evacuation constraints with spatial accessibility theory to produce a robust and interpretable model.
Appendix D Assumptions for Evacuation Congestion and Travel-Speed Reduction
To conservatively represent near-source evacuation congestion under smoke-reduced visibility, we applied a 5 km buffer around evacuation and warning zones, within which vehicular speeds were capped at 10 km/h. This assumption is consistent with wildland–urban interface (WUI) traffic models that permit minimum speeds under dense smoke conditions [36, 63], empirical observations reporting evacuation corridor speeds of approximately 1–3 km/h during recent wildfire events [82], and planning-level analyses that adopt speed thresholds of approximately 5–15 mph (about 8–24 km/h) [21, 3]. In addition, the selected 5 km buffer aligns with the 2–10 km spatial scales commonly used to delineate WUI evacuation buffer zones and study extents in prior research [87, 67, 27].